OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

Well do you know the point slope formula?

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

@halorazer /.\

OpenStudy (anonymous):

No, no. That is slope intercept form. m is the slope and b is the intercept. I am talking about point slope. The equation that, given a point and a slope, you can figure out what the equation is.

OpenStudy (anonymous):

im lost ;-;

OpenStudy (anonymous):

That was a terrible explanation. Point slope is \[y-y1=m(x-x1)\] Slope intercept is \[y=mx+b\] In point slope form, you would put in the coordinates that you were given (2,-5) into their respective spots. (x1 would be 2 and y1 would be -5) then solve from there.

OpenStudy (anonymous):

this doesnt make sense ;-; like nvm sorry for wasting your time ;-;

OpenStudy (anonymous):

i dont understand how i would solve it?

OpenStudy (anonymous):

Okay, well \[y--5=-3(x-2)\] When you subtract a negative number, it switches the sign to a positive, so we get: \[y+5=-3(x-2)\] Can you take it from there?

OpenStudy (anonymous):

would i subtact 2?

OpenStudy (anonymous):

No, first you would distribute the -3 to the x and the -2.

OpenStudy (anonymous):

so its y- -5 +-3- -6?

OpenStudy (anonymous):

Well, you're getting close. Don't forget the equals sign in there. And the x on the -3.

OpenStudy (anonymous):

omg i wrot that on the paper but didnt type it lol sorry cx

OpenStudy (anonymous):

y- -5= -3x- -6

OpenStudy (anonymous):

Yes. Don't forget about the subtracting a negative thing I told you.

OpenStudy (anonymous):

subtract the 3?

OpenStudy (anonymous):

No. y--5 means that it's actually y+5.

OpenStudy (anonymous):

or the 6?

OpenStudy (anonymous):

so i subtact the -6 from the 5 and 3?

OpenStudy (anonymous):

or add?

OpenStudy (anonymous):

I will work through it from the start and explain all my steps.

OpenStudy (anonymous):

okay /.\

OpenStudy (anonymous):

Given the point slope equation \[y-y1=m(x-x1)\] and that we know what the points are (2,-5) we can put them in for y1 and x1 \[y--5=m(x-2)\] When you subtract -5, you're actually adding 5. \[y+5=m(x-2)\] We also know the slope, which is -3. So we can put that in for m. \[y+5=-3(x-2)\] Next we have to distribute the -3 to the x and the -2. \[y+5=-3x+6\] We are trying to get it in slope intercept form which is \[y=mx+b\] Which means that the y has to be completely alone. So in order to get it alone, we have to subtract 5 from both sides of the equation. \[y=-3x+1\] Which is your final option there.

OpenStudy (anonymous):

imma take notes of this /.\ could you help me with 2 small questions? they are litterally small lol

OpenStudy (anonymous):

Which of the following equations best represents the graph of f(n)? f(y) = n f(x) = n x = f(n) y = f(n)

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-8, -7). y = -8 y = -7 x = -8 x = -7

OpenStudy (anonymous):

Is there anything else with the f(n) question?

OpenStudy (anonymous):

no ;-;

OpenStudy (anonymous):

As for the horizontal line, it means that it is staying at a single y coordinate. For example, a horizontal line going through -3,5 would be y=5. So the answer is A.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ill ask some one else , thank you for all the help dear (:

OpenStudy (anonymous):

sorry for bothering you

OpenStudy (anonymous):

I think the answer to the f(n) thing is C. I could be wrong.

OpenStudy (anonymous):

You are most certainly not bothering me. :)

OpenStudy (anonymous):

thank you <3